# DBCopilot: Scaling Natural Language Querying to Massive Databases

### v2

@@ Abstract.
NL2SQL simplifies database interactions by enabling non-experts to convert natural language (NL) questions into Structured Query Language (SQL) queries. While recent advances in large language models (LLMs) have improved the zero/few-shot NL2SQL paradigm, existing methods face scalability challenges when dealing with massive databases. This paper introduces DBCopilot, a framework that addresses these challenges by employing a compact and flexible copilot model for routing over massive databases. Specifically, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation. This framework utilizes a lightweight differentiable search index to construct semantic mappings for massive database schemas and navigates natural language questions to their target databases and tables in a relation-aware, end-to-end manner. The routed schemas and questions are then fed into LLMs for effective SQL generation. Furthermore, DBCopilot introduces a reverse schema-to-question generation paradigm that can automatically learn and adapt the router over massive databases without manual intervention. Experimental results demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, providing a significant advance in handling large-scale schemas of real-world scenarios.
Text-to-SQL, LLM, Generative Retrieval, Schema Routing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Structured Query Languageccs: Information systems Top-k retrieval in databasesccs: Information systems Question answering
Figure 1. The DBCopilot framework. DBCopilot decouples NL querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration.
@@ 1. Introduction
Natural language to SQL (NL2SQL) aims to convert natural language (NL) questions into Structured Query Language (SQL) queries, making it convenient for non-experts to query and analyze databases. This task plays a critical role in data analysis, business intelligence, and other database interaction applications, thus has attracted significant attention (Androutsopoulos et al., 1995; Affolter et al., 2019; Qin et al., 2022; Katsogiannis-Meimarakis and Koutrika, 2023; Fu et al., 2023; Gu et al., 2023a). The recent emergence of large language models (LLMs) brings a new zero/few-shot NL2SQL paradigm that directly transforms SQL queries via LLMs with schema-aware prompts, showing great promise (Rajkumar et al., 2022; Liu et al., 2023a; Pourreza and Rafiei, 2023; Gu et al., 2023b; Gao et al., 2023).
@@
However, existing work in this direction typically focuses only on querying a single database with a handful of tables, neglecting the real-world requirement to query over massive databases and data warehouses (Pride, 2023; Kothyari et al., 2023; Sen et al., 2020). In this case, data consumers need to invest significant effort to manually identify and select the appropriate database and relevant tables. Therefore, it becomes imperative to develop a mechanism that abstracts non-expert users from the intricacies of data representation and enables them to query over massive databases with schema-agnostic NL questions, i.e., without understanding of large-scale schemas. The core of this process is to construct semantic mappings from diverse NL questions to massive database model elements, thus automatically identifying the target database and filtering for a minimal amount of relevant tables.
The above scenarios pose significant challenges to current LLM-based NL2SQL methods. Due to token limits and inference costs, it is infeasible to feed hundreds/thousands of schemas into LLMs for each question. Furthermore, current LLMs still struggle to effectively leverage information present in long contexts (Liu et al., 2023b; Xu et al., 2023). To address this scalability problem, existing solutions can be categorized into retrieval-based and fine-tuning based methods. Given an NL question, retrieval-based methods identify the corresponding table schemas using information retrieval techniques (Wikipedia contributors, 2023a; Karpukhin et al., 2020). Unfortunately, by relying on the co-occurrence of words and representations between question and schema, retrieval-based methods suffer from vocabulary mismatch and the diversity of NL expressions. In addition, these ad hoc retrieval methods typically treat documents as independent objects, neglecting the relationships between databases and tables. Fine-tuning based methods inject schema knowledge into the parameters of LLMs by fine-tuning with schema descriptions or (question, schema) pairs. Unfortunately, fine-tuning LLMs with billions of parameters to keep up with evolving database schemas is resource-intensive and sometimes impractical. Besides, it remains challenging to effectively inject new knowledge into LLMs through fine-tuning (Ovadia et al., 2023). As a result, how to extend the scalability of LLM-based NL2SQL methods and apply them to massive databases is a major challenge in this field.
@@
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. The main idea behind DBCopilot is to introduce a compact and flexible copilot model to serve as a router across massive databases, and to leverage the efficient NL2SQL capability of LLMs to generate final SQL queries. To this end, as shown in Figure 1, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation. Specifically, DBCopilot utilizes a lightweight differentiable search index (DSI) (Tay et al., 2022) to navigate an NL language question to its corresponding database and tables. Then the routed schemas and questions are fed into LLMs to generate SQL queries by leveraging their strong language understanding and SQL generation capabilities. In this way, DBCopilot can easily integrate with advanced LLM-based NL2SQL solutions and rapidly and cost-effectively scale to massive databases.
For an effective implementation of DBCopilot, three major challenges need to be addressed. (1) Schema routing is not a task of retrieving independent tables because complex queries require information across multiple tables, and there are inherent relationships between databases and tables. To resolve this challenge, we model schema routing as a sequence-to-sequence (Seq2Seq) DSI. By proposing schema graph construction, DFS serialization, and constrained decoding, we can jointly retrieve target databases and tables in an end-to-end manner, taking into account structural relationships. (2) The generative retrieval paradigm makes it infeasible to perform semantic mapping of unseen schemas because they are not indexed into model parameters. In practice, labeling NL questions for massive database schemas is highly expensive. To this end, we propose a reverse schema-to-question generation paradigm to automatically synthesize training data and efficiently train the schema router. (3) How to incorporate routed schemas with LLMs to perform schema-aware SQL generation is critical to the quality of NL2SQL. We explore various prompt strategies to select and incorporate multiple candidate schemas for DBCopilot.
@@
We use three standard NL2SQL datasets and two additional robustness datasets to evaluate DBCopilot. Experimental results verify the effectiveness of DBCopilot, with its copilot model outperforming retrieval-based baselines in schema routing by up to 19.88% in recall, and its schema-agnostic NL2SQL showing an improvement in execution accuracy of more than 7.35%. In summary, we demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, presenting a significant step forward in handling large-scale schemas of real-world scenarios.
Contributions.  Our contributions can be summarized as follows.
  • We highlight the limitations of existing NL2SQL methods in scaling from limited tables to massive databases.
@@
  • We present DBCopilot, a framework that decouples schema-agnostic NL2SQL into schema routing and SQL generation via an LLM-copilot collaboration architecture.
@@
  • We propose a relation-aware, end-to-end joint retrieval approach for schema routing, and also propose a reverse generation paradigm for automatic training data synthesis.
@@
  • We conduct thorough experiments to verify the effectiveness of our solution in scaling NL2SQL to massive databases.
@@ 2.1. Problem Definition
Symbol Description
@@
$S$ SQL query schema $\langle D,T\rangle$, where $T$ is the subset of tables belonging to database $D$
@@
$\mathcal{S}$ Schema space $\mathcal{D}\times\mathcal{P}(\mathcal{T})$, where $\mathcal{P}(\mathcal{T})$ is the power set of $\mathcal{T}$
@@
As mentioned above, we formulate the problem of natural language querying over massive databases and large numbers of tables with unknown SQL query schema (i.e., target database and tables) as schema-agnostic NL2SQL. Formally, as per the notation list detailed in Table 1, the problem of schema-agnostic NL2SQL is defined in Definition 2.1, and its two subtasks schema routing and SQL generation are defined in Definition 2.2 and Definition 2.3, respectively. Obviously, traditional NL2SQL with an NL question and its limited database schema as input is similar to the subtask SQL generation.
@@
Given a natural language question $N$, schema-agnostic NL2SQL generates an equivalent SQL query $Q$ that can be executed against the appropriate database $D$ drawn from a collection of databases $\mathcal{D}$, without being provided with the intended SQL query schema $S=\langle D,T\rangle$.
@@
Schema routing is the process of identifying the SQL query schema $S=\langle D,T\rangle$ of a natural language question $N$ from the entire space of all possible schemas $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ within the specified set of databases $\mathcal{D}$ and tables $\mathcal{T}$.
@@
Schema Linking vs. Schema Routing.  Schema linking (Katsogiannis-Meimarakis and Koutrika, 2023) is an optional substep of traditional NL2SQL methods. Given an NL question and a database schema, schema linking aligns NL terminologies, called “mentions”, with their corresponding schema elements (such as tables, columns, or database values). In this way, schema linking servers as a bridge between NL questions and database elements.
Schema linking and schema routing can be distinguished in two aspects. First, schema linking is a substep of SQL generation for accurate generation, while schema routing is a prerequisite of SQL generation to identify the unknown schema of a question. Second, the input for schema linking is the NL question along with the schema, whereas, the input for schema routing is the NL question and the output is its indexed or memorized schema.
@@ 2.2. Generative Retrieval
Traditional information retrieval (IR) techniques typically follow the “index-retrieve” paradigm. However, such a pipelined framework, which cannot be optimized together in an end-to-end manner, limits the final retrieval quality. To address this concern, generative retrieval (Tay et al., 2022; Wang et al., 2022) (a.k.a. DSI) is an emerging paradigm that trains a Seq2Seq pre-trained language model (PLM) to map queries directly to relevant document identifiers. The main idea behind it is to completely parameterize traditional “index-retrieve” pipelines within a single neural model, by modeling indexing and retrieval operations into the model parameters. Critical ongoing research in this paradigm focuses on defining better document identifiers (Sun et al., 2023; Zhang et al., 2023). To the best of our knowledge, this paper is the first to adapt this paradigm to jointly retrieve multiple elements with structural relationships.
@@ 3. Methodology
This section introduces DBCopilot, a schema-agnostic NL2SQL framework for effectively scaling NL querying to massive databases.
@@ 3.1. Overview
As shown in Figure 1, DBCopilot decouples natural language querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration. Specifically, given an NL question $N$, DBCopilot utilizes a copilot model as a schema router to identify the SQL query schema $S=\langle D,T\rangle$ associated with that question. After that, DBCopilot utilizes an LLM to translate the NL question $N$ with the routed schema $S$ into a corresponding SQL query $Q$. Example 3.1 provides an illustration of this process.
@@
Given a question “Which singers held concerts in 2022?”, DBCopilot converts it into an equivalent SQL query through the following two phases.
1. Schema Routing identifies the relevant database and tables that contain the information required to answer the question:
@@
2. SQL Generation prompts the LLM with the identified schema and question, and generates the SQL query:
SELECT s.name FROM singer_in_concert AS sic JOIN singer AS s
ON sic.singer_id = s.singer_id JOIN concert AS c
ON sic.concert_id = c.concert_id WHERE c.year = 2022
Figure 2. The training process of the schema router.
In practice, the schema space $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ is huge because it is the cartesian product of the database collections $\mathcal{D}$ and the power set of all tables $\mathcal{T}$. To narrow down the search space, we propose a relation-aware, end-to-end joint retrieval approach for schema routing. The training process is illustrated in Figure 2. Specifically, we first construct a schema graph to represent the relationships between databases and tables (§ 3.2). Based on this graph, we design a schema serialization algorithm to serialize a SQL query schema into a token sequence so that it can be generated by Seq2Seq DSI (§ 3.3). To address the challenge that DSI cannot generalize to unseen schemas, we further propose a reverse generation paradigm for training data synthesis (§ 3.4). These measures allow us to effectively train the schema router using serialized pseudo-instances ¡$N$, $S$¿ and apply inference with graph-based constrained decoding (§ 3.5). Finally, we explore various prompt strategies to select and incorporate multiple candidate schemas with LLMs (§ 3.6).

### v3

# DBCopilot: Natural Language Querying over Massive Databases via Schema Routing

@@ Abstract.
The development of Natural Language Interfaces to Databases (NLIDBs) has been greatly advanced by the advent of large language models (LLMs), which provide an intuitive way to translate natural language (NL) questions into Structured Query Language (SQL) queries. While significant progress has been made in LLM-based NL2SQL, existing approaches face several challenges in real-world scenarios of natural language querying over massive databases. In this paper, we present DBCopilot, a framework that addresses these challenges by employing a compact and flexible copilot model for routing over massive databases. Specifically, DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation. This framework utilizes a single lightweight differentiable search index to construct semantic mappings for massive database schemata, and navigates natural language questions to their target databases and tables in a relation-aware joint retrieval manner. The routed schemata and questions are then fed into LLMs for effective SQL generation. Furthermore, DBCopilot introduces a reverse schema-to-question generation paradigm that can automatically learn and adapt the router over massive databases without manual intervention. Experimental results verify that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, providing a significant advance in handling natural language querying over massive databases for NLIDBs.
copyright: rightsretaineddoi: isbn: 978-3-89318-099-8conference: 28th International Conference on Extending Database Technology (EDBT); 25th March-28th March, 2025; Barcelona, Spainjournalyear: 2025
@@ 1. Introduction
Natural Language Interfaces to Databases (NLIDBs) enable users to query and interact with databases by using natural language (NL) rather than complex query languages like Structured Query Language (SQL) (Androutsopoulos et al., 1995; Affolter et al., 2019). These systems are critical to data democratization by promoting data accessibility and are widely used in practical scenarios such as data analysis and business intelligence. The development of NLIDBs has been greatly advanced with the advent of large language models (LLMs), which provide an intuitive way to translate NL questions into SQL queries via schema-aware prompts in a new zero- or few-shot NL2SQL paradigm (Rajkumar et al., 2022; Liu et al., 2023; Pourreza and Rafiei, 2023; Gu et al., 2023b; Gao et al., 2024; Li et al., 2024c; Talaei et al., 2024; Li et al., 2024a).
Figure 1. DBCopilot employs LLM-Copilot collaboration to scaling natural language querying over massive databases. Schema routing first navigates to the appropriate database and tables for NL questions via a copilot model. SQL generation then transforms SQL queries via LLMs with schema-aware prompts.
@@
However, existing work in this direction typically focuses only on querying a small database with a handful of tables, neglecting real-world scenarios of NLIDBs to query over massive databases (e.g., data lakes, data warehouses, and open data portals) with large-scale schemata (Pride, 2023; Kothyari et al., 2023; Sen et al., 2020; Stuart, 2015). As illustrated in Figure 1, data consumers in this case need to invest significant effort to manually identify and select the appropriate database and relevant tables. Therefore, it becomes imperative to develop a mechanism that abstracts non-experts from the complex data representation and enables them to query over massive databases with schema-agnostic NL questions, i.e., without understanding of large-scale schemata. The core of this process is to construct semantic mappings from diverse NL questions to massive database model elements, thus automatically identifying the target database and filtering for a minimal amount of relevant tables.
Limitations of state-of-the-art.  The above scenarios pose significant challenges to current LLM-based NL2SQL, schema linking, and ad-hoc retrieval approaches, including:
  1. C1)
    Schema Scalability: It is infeasible to feed hundreds or thousands of schemata into LLMs for each question due to token limits and inference costs. Moreover, current LLMs struggle to effectively leverage information present in long contexts (Liu et al., 2024; Levy et al., 2024). Therefore, numerous studies have employed schema linking as a prerequisite step for LLM-based NL2SQL (Pourreza and Rafiei, 2023; Wang et al., 2025; Li et al., 2024c). However, they follow the typical practice of simultaneously feeding the entire schema and NL question into a model (Dong et al., 2019; Lei et al., 2020; Li et al., 2023c), which hinders their scalability when dealing with large-scale schemata.
  2. C2)
    Schema Complexity: Real-world database schemata are characterized by inherent intricacy and heterogeneity in the relationships between tables, as shown in Figure 1. These relationships are necessary to identify the correct schema because SQL queries for NL questions often span multiple related tables. However, ad-hoc retrieval methods such as BM25 (Wikipedia contributors, 2023b) and DPR (Karpukhin et al., 2020) for schema linking or retrieval-augmented generation (RAG) (Zhao et al., 2024; Gao et al., 2023) merely retrieve elements individually. This approach is suboptimal because the relationships can only be considered through post-processing or re-ranking (Kothyari et al., 2023; Chen et al., 2024) rather than in an end-to-end manner.
  3. C3)
    Semantic Mismatch: Unlike general natural language text, database schemata often employ domain-specific or unconventional terminology and naming conventions for tables and columns (Floratou et al., 2024). This discrepancy entails a semantic mismatch between NL questions and schemata. As a result, retrieval-based methods that rely on the co-occurrence of words or sentence representations suffer from the vocabulary mismatch and the diversity of schemata (Talaei et al., 2024; Li et al., 2024c). While some efforts attempt to rewrite NL questions into possible schema elements via LLMs for format alignment (Kothyari et al., 2023; Chen et al., 2024), they do not adequately address the underlying semantic mismatch.
@@ 3. C3)
Approach.  In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. The main idea behind DBCopilot is to introduce a compact and flexible copilot model for routing over massive databases, and to leverage a powerful LLM for generating SQL queries. By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot avoids frequently adapting LLMs to domain-specific data while maintaining their generic SQL generation capability. As shown in Figure 1, DBCopilot utilizes a lightweight Differentiable Search Index (DSI) (Tay et al., 2022) to navigate an NL question to its corresponding database and tables. Then the routed schemata and questions are fed into LLMs to generate SQL queries by leveraging their strong capabilities. In this way, DBCopilot can be easily integrated with advanced LLM-based NL2SQL solutions and scale rapidly and cost-effectively to massive databases for NLIDBs.
To effectively implement DBCopilot, we address the above challenges in the following way. For C1) schema scalability, we leverage generative retrieval (a.k.a. DSI) to avoid the need to simultaneously feed whole schemata and the NL question into a neural model. By injecting schema knowledge into the parameters of the schema router, the corresponding SQL query schema can be generated by simply inputting an NL question, thereby extending scalability. For C2) schema complexity, we propose a relation-aware joint retrieval approach to identify the target databases and tables in an end-to-end manner. We first construct a schema graph to represent the relationships between databases and tables. Based on this graph, we design DFS serialization and constrained decoding algorithms to facilitate relation-aware Seq2Seq retrieval. For C3) semantic mismatch, we introduce a reverse schema-to-question generation paradigm to automatically synthesize mass training data and efficiently learn the semantic mapping for the schema router. This approach effectively mitigates the labeling cost for massive database schemata and the problem that generative retrieval cannot generalize to unseen schemata during training. Finally, we explore various prompt strategies to select and incorporate multiple candidate schemata for LLM-based SQL generation.
@@
Evaluation.  We adapt three standard NL2SQL datasets and two robustness datasets to evaluate DBCopilot. Experimental results verify the effectiveness of DBCopilot, with its copilot model outperforming retrieval-based baselines in schema routing by up to 19.88% in recall, and its schema-agnostic NL2SQL showing an improvement in execution accuracy of 4.43%~11.22%. In summary, we demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, presenting a significant step forward in handling large-scale schemata of real-world scenarios.
Contributions.  Our contributions are summarized as follows. [^1] We have open-sourced our code and data at: github.com/tshu-w/DBCopilot.
  • We approach the practical and challenging problem of NLIDBs for querying over massive databases, while pointing out the limitations of existing approaches.
@@
  • We present DBCopilot, a framework that addresses this problem by decoupling schema-agnostic NL2SQL into schema routing and SQL generation via LLM-copilot collaboration.
@@
  • We propose a relation-aware joint retrieval approach for end-to-end schema routing and a reverse generation paradigm for learning semantic mappings via training data synthesis.
@@
  • We conduct thorough experiments to verify the effectiveness of DBCopilot in scaling NL2SQL to massive databases.
@@ 2.1. Problem Formulation
[…]
notes/style=0 {NiceTabularX} 1.13cX Symbol Description
@@
$S$ SQL query schema $\langle D,T\rangle$, where $T$ is the subset of tables in database $D$
@@
[…]
@@
As mentioned above, we abstract and formulate the problem of natural language querying over massive databases and large numbers of tables with unspecified SQL query schema (i.e., the target database and tables) as schema-agnostic NL2SQL. Formally, as per the notation list detailed in § 2.1, the problem of schema-agnostic NL2SQL and its two subtasks schema routing and SQL generation are described in Definition 2.1, 2.2, and 2.3, respectively.
@@
Given a natural language question $N$, schema-agnostic NL2SQL generates an equivalent SQL query $Q$ that can be executed against the appropriate database $D$ drawn from a set of databases $\mathcal{D}$, without being provided with the intended SQL query schema $S=\langle D,T\rangle$.
@@
Schema routing identifies the SQL query schema $S=\langle D,T\rangle$ of a natural language question $N$ from the space of all possible schemata $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ within the specified set of databases $\mathcal{D}$ and tables $\mathcal{T}$.
@@
Schema Linking vs. Schema Routing.  Schema linking is an optional substep of NL2SQL methods (Katsogiannis-Meimarakis and Koutrika, 2023). Given an NL question and a database schema, schema linking aligns natural language terminologies, called “mentions”, with their corresponding schema elements (such as tables, columns, and database values). In this way, schema linking servers as a bridge between NL questions and database elements.
Schema routing is an extension of the schema linking concept adapted for querying over massive databases. Specifically, schema routing differs from existing schema linking methods in the following aspects. First, schema linking methods (Dong et al., 2019; Lei et al., 2020; Li et al., 2023c) usually take the entire single-database schema and an NL question as input, while schema routing, driven by scalability requirements, should take the NL question as input with massive database schemata indexed. Second, schema routing care about coarse-grained databases and tables, while leaving the fine-grained columns and values for further SQL generation.
@@ 2.2. Generative Retrieval
Traditional information retrieval (IR) techniques follow the “index-retrieve” paradigm. Such a pipeline, which cannot be optimized together in an end-to-end manner, limits the final retrieval quality. To address this concern, generative retrieval (Tay et al., 2022; Wang et al., 2022) (a.k.a. DSI) is an emerging paradigm that leverages a Seq2Seq pre-trained language model (PLM) to directly generate relevant document IDs (e.g., “doc456”) in response to input queries, rather than vector representations for computing similarities. Technically, DSI works by Seq2Seq learning on (query/doc_chunk, doc_id) pairs to model semantic mappings, completely parameterizing traditional “index-retrieve” pipelines within a single neural model. Therefore, unlike conventional machine learning tasks that aim to generalize to unseen data, the training phase of DSI shifts to memorize all retrieval targets – including those used in evaluation – similar to the indexing phase of traditional retrieval. Critical ongoing research focuses on defining better document IDs (Sun et al., 2023; Zhang et al., 2023). However, these studies mainly focus on retrieving elements individually without considering their relationships (e.g., joinable relations between tables, triple relations in knowledge graphs, or semantic relations between documents). To the best of our knowledge, this paper is the first to adapt this paradigm to jointly retrieve elements with relationships, opening up new opportunities for applying it in graph-based retrieval (Peng et al., 2024).
2.3. LLM-based Natural Language to SQL
LLMs, developed by scaling up PLMs in terms of both model and data size, have emerged as a milestone in artificial intelligence (Zhao et al., 2023). Despite their similar architecture and pre-training tasks, LLMs exhibit unexpected emergent abilities (e.g., in-context reasoning, instruction following) compared to small PLMs (Wei et al., 2022). These emergent abilities change the paradigm of model use from “fine-tuning” to “prompt engineering” that does not change the model parameters. This shift has brought new opportunities for all kinds of natural language processing tasks, including NL2SQL. Coupled with well-designed schema-aware prompts and few-shot NL2SQL examples, LLMs can generate SQL queries directly from NL questions, delivering state-of-the-art performance (Gao et al., 2024; Gu et al., 2023b; Li et al., 2024c). In addition, pipeline-based approaches decompose NL2SQL into substeps including schema linking, query classification, query decomposition, SQL refinement, further extending the power of LLM-based NL2SQL (Pourreza and Rafiei, 2023; Li et al., 2024b). Nevertheless, LLM-based NL2SQL still faces significant challenges in real-world scenarios, such as schema scalability, complex or ambiguous queries, SQL efficiency, and so on (Floratou et al., 2024; Li et al., 2023b; Kothyari et al., 2023).
@@ 3. Methodology
This section introduces DBCopilot, a schema-agnostic NL2SQL framework for effectively scaling natural language querying to massive databases.
Figure 2. Training process of schema router. We first construct a schema graph to represent the relationships of all databases and their tables. Based on this graph, we sample SQL query schemata and generate pseduo-questions using a reverse-trained schema questioning model that synthesizes training data. Finally, we apply DFS serialization and train the schema router.
@@ 3.1. Overview
As shown in Figure 1, DBCopilot decouples natural language querying over massive databases into schema routing and SQL generation. Specifically, given an NL question $N$, DBCopilot utilizes a copilot model as a schema router to identify the SQL query schema $S=\langle D,T\rangle$ associated with that question. After that, DBCopilot utilizes an LLM to translate the NL question $N$ with the routed schema $S$ into a corresponding SQL query $Q$. Example 3.1 provides an illustration of this process. Consequently, this LLM-Copilot collaboration architecture enables easy integration with advanced LLM-based NL2SQL solutions and cost-effective scaling to massive databases for NLIDBs.
@@
Given a natural language question “Which singers held concerts in 2022?”, DBCopilot converts it into an equivalent SQL query through the following two phases:
  1. 1.
    Schema Routing identifies the relevant database and tables containing necessary information to answer the question.
@@
  2. 2.
    SQL Generation prompts LLMs with the identified schema and NL question to generate the SQL query.
    SELECT s.NAME
    FROM concert_singer.singer_in_concert AS sic
    JOIN concert_singer.singer AS s
    ON sic.singer_id = s.singer_id
    JOIN concert_singer.concert AS c
    ON sic.concert_id = c.concert_id
    WHERE c.year = 2022
To better scale schema routing to massive databases by modeling relationships between schema elements and learning semantic mappings from natural language to schemata, we propose a relation-aware joint retrieval approach for end-to-end schema routing. Specifically, we first construct a schema graph to represent the underlying relationships between databases and tables (§ 3.2). Based on this graph, we design a schema serialization algorithm to serialize a SQL query schema into a token sequence so that it can be generated by Seq2Seq DSI (§ 3.3). To bridge the semantic gap between natural language and domain-specific schemata, we further propose a reverse generation paradigm for training data synthesis (§ 3.4). These measures allow us to effectively train the schema router with serialized pseudo-instances <$N$, $S$> and apply inference with graph-based constrained decoding (§ 3.5). Figure 2 illustrates the training process. Finally, we explore various prompt strategies to select and incorporate multiple candidate schemata for LLM-based SQL generation (§ 3.6).